Graph to solve the following linear system.
Y=-x+5
-5x+2y=10

Answer by mananth(16946)

Y=-x+5.......................(1)
-5x+2y=10.....................(2)
Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= 5 ,( 0 , 5 )
when y = 0 x= 5 ,( 5 , 0 )
Equation 2
when x= 0 y= 5 ( 0 , 5 )
when y = 0 x= -2 ( -2 , 0 )



You will observe that the point of intersection is 0 , 5
The solution of the system is ( 0 , 5 )